OpenStudy (anonymous):

solve for positive exponents -20x^3y^4 80x^2y^-1

OpenStudy (anonymous):

is there division...?

OpenStudy (anonymous):

there is a line in between

OpenStudy (anonymous):

-20/80 = 1/4 x^3/x^2 = x y^4/y^-1 = y^5 so, your answer should be: (-1/4) *x*y^5